# Optimal quantum transport on a ring via locally monitored chiral quantum walks

### v1

@@ Abstract
I Introduction
In a classical random walk the walker takes steps in random directions through a network, resulting in diffusive spreading. In a quantum walk, instead, the walker takes a quantum superposition of paths, leading to quantum interference effects [1, 2]. While constructive interference enables ballistic spreading (faster than the classical diffusive spreading), destructive interference can hinder transport or cause localization [3].
@@ I Introduction
Continuous-time quantum walks are versatile for modeling coherent transport in discrete systems [4]. Equivalent to the tight-binding approximation in solid-state physics and in Hückel’s molecular-orbital theory, their applicability is broad across systems, ranging from solid-state to biological ones [5, 6, 7, 8, 9, 10, 11]. In recent decades, several studies on environment-assisted quantum transport have highlighted the beneficial interplay of coherent and incoherent dynamics [12, 13, 14, 15, 16], where the loss of phase coherence, e.g., due to dephasing, can enhance transport by suppressing destructive interference [17, 18]. Focusing on purely coherent transport, achieving unit transport efficiency in highly symmetric networks often requires non-trivial delocalized initial states due to symmetry-protected states [19, 20]. Breaking some symmetries, e.g., via network defects [21, 22], can improve transport.
@@
Chiral quantum walks leverage time-reversal symmetry breaking, typically through complex-valued edge weights [23, 24], to enable controlled directional transport [25, 26, 27] or its suppression [28], with promising applications in routing [29]. While chirality can mitigate destructive interference by affecting phase coherences, achieving nearly optimal transfer—even on a simple ring [26]—often requires fine control of the evolution time [30]. This represents an experimental challenge, as it requires a priori knowledge of the optimal time to halt the transfer protocol. Classically, this relates to first-passage-time problems [31, 32, 33, 34], which address the statistics of the time a random walker takes to reach a target. However, in a quantum context the concept of first passage is not meaningful; instead, measurements must be explicitly incorporated [19], leading to the notion of first-detected-passage time [35, 36, 37, 38, 39]. The latter bridges unitary dynamics and measurement-induced collapse in monitored systems [40], offering insights into quantum-classical transition and measurement back-action.
@@
While it is known that chirality can mitigate destructive interference by lifting energy degeneracies, associated to dark states that prevent detection for certain initial conditions [38, 41, 42], the optimal interplay between chirality and measurement frequency has yet to be explored. In this work, we address this gap by investigating a monitored chiral quantum walk on a ring (see Fig. 1), where the purely coherent evolution of the system is repeatedly interrupted by projective measurements at the target. We show how to leverage chirality and detection period to significantly enhance transport. Specifically, we derive a general prescription for maximizing the asymptotic detection probability, relying on the Perron-Frobenius spectrum of the non-unitary dynamics and on the analysis of dark states.
The paper is organized as follows: In Section II, we introduce the chiral system under investigation and briefly review the first-detected-passage time problem. In Section III we present results about the maximization of the detection probability. Relying on the analysis of the Perron-Frobenius spectrum and dark states, in Section IV we discuss the optimal conditions for excitation transfer in the long-time limit and draw conclusions in Section V.
@@
II First-detected-passage-time problem
Within the framework of first-detected-passage-time problems, our purpose is to investigate coherent excitation transfer on a ring—a cycle graph with finite number of sites $N$—under local monitoring. We model the transfer as a continuous-time chiral quantum walk, and assume stroboscopic detection for local monitoring (see Fig. 1). For this system, a convenient basis of the Hilbert space is represented by the localized states on the sites of the graph, {|$\{\ket{j}\}_{j=0,\dots,N-1}$. In such a basis, the Hamiltonian of the chiral quantum walk can be written as [25]
@@ II First-detected-passage-time problem
|$\ket{\lambda_{j}}$ $=\frac{1}{\sqrt{N}}\sum_{k=0}^{N-1}e^{i\frac{2\pi}{N}jk}\ket{k},$ (2)
$\lambda_{j}$ $=2\cos\left(\phi-\frac{2\pi}{N}j\right),$ (3)
@@
In quantum systems the first-detected-passage time, or hitting time, is defined through repeated monitoring at the target site [19, 37], where the detector reveals the presence or absence of the walker. We assume a stroboscopic detection protocol in which projective measurements are performed at times $\tau,2\tau,3\tau,\ldots$, where the arbitrary finite sampling period $\tau>0$ is chosen by the experimentalist. The scheme is illustrated in the box in Fig. 1: starting from an initial state |$\ket{\psi_{0}}=\ket{0}$ localized at site (without loss of generality we assume the 0th), the monitored dynamics alternates unitary evolution, $\hat{U}(\tau)$ (4), and projective measurements at the target site, $\hat{D}=\outerproduct{\delta}{\delta}$ |. In the following, we assume the target site to be the opposite to the initial one, i.e., |$\ket{\delta}=\ket{N/2}$ for even $N$, and |$\ket{\delta}=\ket{(N\pm 1)/2}$ for odd $N$. A typical run of the monitored dynamics produces a string of binary measurement outcomes: a sequence of no (the walker has not been detected) repeated until a yes (the walker has been detected) occurs at time $n\tau$, i.e., at the $n$th detection attempt. This time $n\tau$ is then defined as the first-detected-passage time for the run under investigation. The probability $F_{n}$ to first detect the walker at the $n$th attempt is [37]
$F_{n}=\bra{\theta_{n}}\hat{D}\ket{\theta_{n}},$ (5)
where
|$\ket{\theta_{n}}=U(\tau)[(\mathbb{I}-\hat{D})U(\tau)]^{n-1}\ket{\psi_{0}},$ (6)
is the first-detection (unnormalized) vector. The detection probability up to time $n\tau$ is
@@
As a final remark, we point out that we investigate local monitoring for excitation transfer since purely-coherent transport performs poorly. Indeed, under unitary dynamics, the transfer probability $P_{\delta}(t)=|\langle\delta|\hat{U}(t)|\psi_{0}\rangle|^{2}$ typically exhibits low values with rare sharp peaks (see Appendix A). By leveraging chirality, $\phi$, and local monitoring, $\tau$, we aim at finding a robust optimal transfer protocol, meaning that no fine-tuning of parameters is required to achieve sufficiently high values of detection probability.
III Maximization of the detection probability
@@
In this section, we aim at maximizing the detection probability (7) over the key parameters of our model: the detection period $\tau$ between consecutive measurements, and the phase $\phi$ in the Hamiltonian (1). This optimization is subject to a constraint, as we assume the total time $T$ of the process to be a finite resource. Therefore, what we maximize is the detection probability $P_{\rm det}(\phi,\tau)$ after $n$ detection attempts, where $n=\lfloor T/\tau\rfloor$. Therefore, the detection probability we discuss is upper bounded by its asymptotic value, $P_{\rm det}(\phi,\tau)\leq P_{\rm det}(n\to\infty)$. In the following, we conveniently set $T=200$, a value which allows us to achieve detection probabilities $>90\%$ when the number of sites $N$ is relatively small (see Section IV.3), while keeping the number of measurements $n$ limited. In the next sections, we will focus first on the role of the phase $\phi$, and then on that of the period $\tau$.
@@
Figure 2: Detection probability as a function of the phase $\phi$ and the sampling period $\tau$ at fixed total observation time $T=200$. Density plots $P_{\rm det}(\phi,\tau)$ (a) for $N=20$ and $\delta=N/2$ and (b) for $N=21$ and $\delta=(N-1)/2$. (c,d) Curves $P_{\rm det}(\phi)$ at given $\tau$ for system size $N$ and detection site $\delta$ as in (a,b), respectively.
III.1 Optimal phase $\phi$
@@
To get a first insight on how the detection probability depends on the parameters of interest, in Fig. 2 we show the density plots of $P_{\rm det}(\phi,\tau)$ for a fixed number of sites, distinguishing between even $N=20$ and odd $N=21$. For even $N$ [Fig. 2(a)], we observe that for $\tau<\tau^{*}$ the detection probability is maximized at $\phi=0$, decreases as $|\phi|$ increases, and vanishes at $\phi=\pm\pi/N$. Above this threshold $\tau^{*}\approx 1.55$ on the detection interval (see Section IV.2), complex patterns emerge. Therefore, introducing a nonzero phase in the Hamiltonian (chirality) hinders excitation transfer between opposite sites of the symmetric cycle (even $N$), as it reduces the detection probability at the target. In contrast, chirality enhances this transfer in the asymmetric cycle (odd $N$) [Fig. 2(b)]. The directionality induced by the nonzero phases causes the detection probability to increase at either of the two target sites $\delta=(N\pm 1)/2$ opposite to the starting one. In particular, for $\tau<\tau^{*}$ the detection probability is always maximized at $\phi_{\rm opt}=\pm\pi/2N$, and minimized at $\phi=0,\pm\pi/N$. As a final remark, we point that while $P_{\rm det}(\phi,\tau)=P_{\rm det}(-\phi,\tau)$ for even $N$, this symmetry is lifted for odd $N$ [Fig. 2(c,d)]. However, for odd $N$ the detection probability is symmetric upon changing both the target site and the sign of $\phi$.
III.2 Optimal measurement period $\tau$
@@
The detection period $\tau$ plays a crucial role, as the detection probability for both even and odd $N$ shows two qualitatively distinct regimes separated by a sharp transition at a critical threshold $\tau^{*}$ [see Fig. 2(a,b)]. When $\tau<\tau^{*}$, $P_{\rm det}(\phi,\tau)$ is smooth, generally increasing with $\tau$ (recall we fixed $T=n\tau$), and remains high over extended intervals. Conversely, when $\tau>\tau^{*}$, a highly structured pattern with sharp oscillations emerges. As $\tau\to 0$, we observe a quantum Zeno effect [44, 45]. Here, frequent measurements performed on a state |$\ket{\delta}$ (orthogonal to the initial one) effectively confine the walker’s evolution in the subspace orthogonal to the detection subspace, preventing it from reaching |$\ket{\delta}$. As a result, the detection probability approaches zero as $\tau\to 0$. A qualitatively analogous behavior is observed when varying the number of sites $N$, with the critical threshold $\tau^{*}$ exhibiting only minor variation. This is illustrated in Fig. 3, where $P_{\rm det}(\phi_{\rm opt},\tau)$ is plotted for different $N$. As evident from Figs. 2-3, although maxima of the detection probability also occur in the region $\tau>\tau^{*}$, attaining them would require a fine tuning of $\tau$ that is difficult to achieve in practice. Since we are interested in a robust protocol, we maximize the detection probability in the relatively wide region $\tau<\tau^{*}$.
@@
IV Analysis and discussion
In this section, we offer two complementary interpretations, one based on the analysis of the Perron-Frobenius spectrum and the other on that of dark states, of the results presented in Section III concerning the optimal parameters $\phi_{\rm opt}$ and $\tau_{\rm opt}$, as well as the presence of the threshold $\tau^{*}$.
IV.1 Perron-Frobenius analysis
Local monitoring repeatedly interrupts the unitary evolution of the system by projective measurements, resulting in an overall non-unitary evolution. The elementary step—a free evolution (4) for time $\tau$ followed by the detection attempt $\hat{D}=\outerproduct{\delta}{\delta}$ |—is governed by the (non-unitary) Perron–Frobenius operator [^1] The authors of Ref. [38] refer to this operator as the survival operator.
@@ IV.1 Perron-Frobenius analysis
$S(n)$ $=\sum_{i=0}^{N-1}|\langle i|\hat{O}^{n}(\tau)|\psi_{0}\rangle|^{2}$
$=\sum_{i=0}^{N-1}\sum_{j=0}^{N-1}|\mu_{j}|^{2n}|\langle i|\mu_{j}\rangle|^{2}|\langle\mu_{j}|\psi_{0}\rangle|^{2},$ (9)
@@
is the probability that the detector has not clicked after $n$ detection attempts, meaning that the walker has not been detected yet [38]. The survival probability relates to the detection probability (7) via $S(n)=1-P_{\rm det}(n)$. The fact that $S(n)\leq 1$ means that $|\mu_{j}|\leq 1\;\forall j$. In particular, the contributions from eigenvectors with $|\mu_{j}|<1$ vanish for $n\to\infty$, while those from eigenvectors with $|\mu_{j}|=1$ survive. The asymptotic dynamics is thus determined by the largest-modulus eigenvalues and their corresponding eigenvectors. The analysis of the eigenproblem of $\hat{O}(\phi,\tau)$ will equip us with the necessary tools to understand both the presence of the threshold $\tau^{*}$ and the optimal values of the parameters. For clarity of discussion, we focus here on the case of odd $N$, deferring the case of even $N$ to Appendix B.
@@
First, we focus on the existence of the threshold $\tau^{*}$ by studying how the Perron-Frobenius spectrum—evaluated at fixed $N$ and $\phi=\phi_{\rm opt}$—differ depending on the regime $\tau<\tau^{*}$ or $\tau>\tau^{*}$ [Fig. 4(a,b)]. For $0<\tau<\tau^{*}$, all the eigenvalues fall within the unit circle, $|\mu_{j}|<1\;\forall j$ [see also the largest-modulus eigenvalue in Fig. 4(d)], meaning that all the modes decay and so the excitation will be transferred to the target state with unit probability in the long-time limit, $P_{\rm det}(n\to\infty)=1$. For $\tau>\tau^{*}$, instead, there may exist unit-modulus eigenvalues, indicating the emergence of modes which do not decay over time. This finding based on the analysis of the Perron-Frobenius spectrum aligns with the observations made in the previous sections, providing an estimate of $\tau^{*}$ that is consistent with the numerically observed one, $\tau^{*}\approx 1.55$. In turn, the value $\tau^{*}$ represents the upper bound to the interval between detection, over which maximizing the detection probability  [^2] Note that, regardless of $\tau$, $\phi$, and $N$ (even or odd), we always observe the presence of a null eigenvalue with corresponding eigenstate $|\mu=0\rangle=\hat{U}(\phi,-\tau)|\delta\rangle$, which trivially leads to $\hat{O}(\phi,\tau)|\mu=0\rangle\equiv 0$..
@@
The optimal parameters, $\tau_{\rm opt}$ and $\phi_{\rm opt}$, can be estimated by minimizing the Perron-Frobenius eigenvalue $\mu_{\rm PF}$ with the largest modulus, which governs the asymptotic dynamics in the long-time limit. The study of $|\mu_{\rm PF}|$ as a function of $\phi$ and $\tau$ in Fig. 4(c) reveals the following. At $\phi=\phi_{\rm opt}=\pm\pi/2N$ (odd $N$), $|\mu_{\rm PF}|$ decreases smoothly with increasing $\tau<\tau^{*}$, attains a minimum at $\tau_{\rm opt}$, and for $\tau>\tau^{*}$ its behavior becomes highly oscillatory, with $|\mu_{\rm PF}|=1$ for specific values of $\tau$ [Fig. 4(d)]. At a given measurement period $\tau<\tau^{*}$, $|\mu_{\rm PF}|$ is minimum at the optimal phase $\phi_{\rm opt}$ [Fig. 4(e)]. This behavior with respect to $\phi$ and $\tau$ is consistent with that of the detection probability in Fig. 2(b,d).
@@
A remark is in order. The Perron-Frobenius analysis characterizes the asymptotic behavior in the long-time limit, whereas our previous investigation (Section III) assumes a finite observation time $T$, which may not correspond to the long-time limit (we will discuss it in Section IV.3). In this regard, letting $\tau_{\rm PF}$ be the value for which $|\mu_{\rm PF}|$ is minimum, we observe that $\tau_{\rm PF}$ is well approximated by $\tau_{\rm opt}$ for low $N$, and this agreement extends to higher $N$ upon increasing the total observation time $T$ (data not shown).
These results show that the Perron-Frobenius analysis can be used to determine the optimal measurement period, provided the total observation time is large enough. As a final remark, we point out that $\tau_{\rm PF}$ approaches the value $\pi/2$ as $N$ increases, a behavior that we will explain in terms of dark states in the next Section IV.2  [^3] Note that this trend is not captured by Fig. 3, where the maximum of the detection probability shifts instead toward lower $\tau$, due the finite observation time $T$..
@@
Figure 4: Eigenvalues of the Perron-Frobenius operator (IV.1) on the unit circle for two representative values of the sampling period, (a) $\tau<\tau^{*}$ and (b) $\tau>\tau^{*}$ at $\phi=\pi/2N$, $\delta=10$. (c) Density plot of the largest-modulus PF eigenvalue $|\mu_{\rm PF}|$ as a function of $\phi$ and $\tau$. (d) Curve $|\mu_{\rm PF}|$ as a function of $\tau$ at fixed $\phi=\pi/2N$. (e) Curve $|\mu_{\rm PF}|$ as a function of $\phi$ at fixed $\tau=1.53$. Other parameters are: $N=21$, $\delta=10$.
IV.2 Dark states
@@
A complementary explanation of the optimal value of the measurement period can be provided using the concept of dark states. As thoroughly explained in [38], the presence of a detector in a finite-dimensional quantum system, such as ours, splits the Hilbert space into two parts: a bright subspace, whose states are guaranteed to be eventually detected, ($P_{\rm det}(\infty)=1$), and a dark subspace, whose states are never detected, not even after an infinite number of detection attempts ($P_{\rm det}(\infty)=0$). A state that lies in the dark subspace is orthogonal (and remains orthogonal under the action of the evolution operator) to the detection site $|\delta\rangle$. A state that satisfies these properties—i.e., it is stationary with respect to the evolution and detection attempts—is called a dark state.
The total detection probability $P_{\rm det}(\infty)$ is the squared modulus of the projection of the initial state $|\psi_{0}\rangle$ on the bright space [38]. Accordingly, whenever $|\psi_{0}\rangle$ has a nonzero projection on the dark space, then $P_{\rm det}(\infty)<1$, meaning that we are not guaranteed to detect the walker even with infinite measurements. Understanding when dark states arise and characterizing them is therefore important to optimize the detection probability.
@@
Considering the time evolution in Eq. (4), a dark state can be built using the eigenvectors (2) of the Hamiltonian (1),

### v2

@@ Abstract
Introduction.—In a classical random walk the walker takes steps in random directions through a network, resulting in diffusive spreading. In a quantum walk, instead, the walker takes a quantum superposition of paths, leading to quantum interference effects [1, 2]. While constructive interference enables ballistic spreading (faster than the classical diffusive spreading), destructive interference can hinder transport or cause localization [3]. Continuous-time quantum walks are versatile for modeling coherent transport in discrete systems [4]. Equivalent to the tight-binding approximation in solid-state physics and in Hückel’s molecular-orbital theory, their applicability is broad across systems [5, 6, 7, 8, 9, 10, 11]. Several studies on environment-assisted quantum transport have highlighted the beneficial interplay of coherent and incoherent dynamics [12, 13, 14, 15, 16], where the loss of phase coherence can enhance transport by suppressing destructive interference [17, 18]. Focusing on purely coherent transport, achieving unit transport efficiency in highly symmetric networks often requires non-trivial delocalized initial states due to symmetry-protected states [19, 20]. Breaking some symmetries, e.g., via network defects [21, 22], can improve transport.
@@
Chiral quantum walks leverage time-reversal symmetry breaking, typically through complex-valued edge weights [23, 24], to enable controlled directional transport [25, 26, 27] or its suppression [28], with promising applications in routing [29]. While chirality can mitigate destructive interference by affecting phase coherences, achieving nearly optimal transfer—even on a simple ring [26]—often requires fine control of the evolution time [30]. This represents an experimental challenge, as it requires a priori knowledge of the optimal time to halt the transfer protocol. Classically, this relates to first-passage-time problems [31, 32, 33, 34], which address the statistics of the time a random walker takes to reach a target. However, in a quantum context the concept of first passage is not meaningful; instead, measurements must be explicitly incorporated [19], leading to the notion of first-detected-passage time [35, 36, 37, 38, 39]. The latter bridges unitary dynamics and measurement-induced collapse in monitored systems [40], offering insights into quantum-classical transition and measurement back-action. Although chirality has been shown to enhance transfer probabilities by lifting energy degeneracies that give rise to dark states—initial conditions which, suffering from destructive interference, prevent the detection of the desired state [38, 41, 42]—the optimal interplay between chirality and measurement frequency remains unexplored.
@@
In this Letter, we address this gap by investigating a locally monitored chiral quantum walk on a ring (see Fig. 1), where the purely coherent evolution of the system is repeatedly interrupted by projective measurements at the target. First, we show the effective role of chirality and detection period in enhancing the detection probability at finite time. Then, we identify the optimal conditions for maximizing the (asymptotic) detection probability, deriving a general prescription that relies on the Perron-Frobenius spectrum of the non-unitary dynamics and on the analysis of dark states. Finally, we elaborate on the time scale for the asymptotic dynamics.
@@
Model and first-detected-passage-time problem.—We investigate coherent excitation transfer on a cycle graph with $N$ sites (a ring) under local monitoring (see Fig. 1). The transfer is modeled as a continuous-time chiral quantum walk and the Hilbert space of the system is conveniently spanned by localized states at the sites of the graph, {|$\{\ket{j}\}_{j=0,\dots,N-1}$. In this basis, the Hamiltonian of the chiral quantum walk can be written as [25]
@@
|$\ket{\lambda_{j}}=\frac{1}{\sqrt{N}}\sum_{k=0}^{N-1}e^{i\frac{2\pi}{N}jk}\ket{k},\quad\lambda_{j}=2\cos\bigg(\phi-\frac{2\pi}{N}j\bigg{missing}),$ (2)
@@
In quantum systems the first-detected-passage time, or hitting time, is defined through repeated monitoring at the target site [19, 37], where the detector reveals the presence or absence of the walker. Here, we assume a stroboscopic detection protocol in which projective measurements are performed at times $\tau,2\tau,3\tau,\ldots$, where the arbitrary finite detection period $\tau>0$ is chosen by the experimentalist. The scheme is illustrated in the box in Fig. 1: starting from an initial state |$\ket{\psi_{0}}=\ket{0}$ localized at a site (without loss of generality we assume the 0th), the monitored dynamics alternates unitary evolution, $\hat{U}(\tau)$ (3), and projective measurements at the target site, $\hat{D}=\outerproduct{\delta}{\delta}$ |. In the following, we assume the target site to be the opposite to the initial one, i.e., |$\ket{\delta}=\ket{N/2}$ for even $N$, and |$\ket{\delta}=\ket{(N\pm 1)/2}$ for odd $N$. A typical run of the monitored dynamics produces a string of binary measurement outcomes: a sequence of no (the walker has not been detected) repeated until a yes (the walker has been detected) occurs at time $n\tau$, i.e., at the $n$th detection attempt. This time $n\tau$ is then defined as the first-detected-passage time for the run under investigation. The probability $F_{n}$ to first detect the walker at the $n$th attempt is $F_{n}=\bra{\theta_{n}}\hat{D}\ket{\theta_{n}}$, where |$\ket{\theta_{n}}=U(\tau)[(\mathbb{I}-\hat{D})U(\tau)]^{n-1}\ket{\psi_{0}}$ is the first-detection (unnormalized) vector [37]. The detection probability up to time $n\tau$ is
@@
As a final remark, we point out that we investigate local monitoring for excitation transfer since purely-coherent transport performs poorly. Indeed, under unitary dynamics, the transfer probability $P_{\delta}(t)=|\langle\delta|\hat{U}(t)|\psi_{0}\rangle|^{2}$ typically exhibits low values with rare sharp peaks (see Sec. S1 of Supplemental Material (SM) [44]).
@@
Optimization problem.—Our purpose is to leverage the key parameters of our model—the detection period $\tau$ between consecutive measurements, and the phase $\phi$ in the Hamiltonian (1)—to determine an optimal, robust transfer protocol: optimal, as it maximizes the detection probability (4) over $\phi$ and $\tau$, and robust, as it does not require fine-tuning of parameters. This optimization is subject to a constraint, as we assume the total time $T$ of the process to be a finite resource. Therefore, what we maximize is the detection probability $P_{\rm det}(\phi,\tau)$ after $n$ detection attempts, where $n=\lfloor T/\tau\rfloor$, which is clearly upper bounded by its asymptotic value, $P_{\rm det}(\phi,\tau)\leq P_{\rm det}(n\to\infty)$. In the following, we conveniently set $T=200$, a value which allows us to achieve detection probabilities $>90\%$ when the number of sites $N$ is relatively small (see the discussion on the asymptotic time scale at the end of the Letter), while keeping the number of measurements $n$ limited. Hereafter, we will focus first on the role of the phase $\phi$, and then on that of the detection period $\tau$.
@@
Figure 2: Detection probability as a function of the phase $\phi$ and the detection period $\tau$ at fixed total observation time $T=200$. Density plots $P_{\rm det}(\phi,\tau)$ (a) for $N=20$ and $\delta=N/2$ and (b) for $N=21$ and $\delta=(N-1)/2$. (c,d) Curves $P_{\rm det}(\phi)$ at given $\tau$ for system size $N$ and detection site $\delta$ as in (a,b), respectively.
@@
Optimal phase $\phi$.—To gain initial insight into how the detection probability depends on the parameters of interest, Fig. 2 shows the density plots of $P_{\rm det}(\phi,\tau)$ for a fixed number of sites, comparing even ($N=20$) and odd ($N=21$) cases. For even $N$ [Fig. 2(a)], we observe that for $\tau<\tau^{*}$ the detection probability is maximized at $\phi=0$, decreases as $|\phi|$ increases, and vanishes at $\phi=\pm\pi/N$. Above this threshold $\tau^{*}\approx 1.55$ in the detection period, complex patterns emerge. Therefore, introducing a nonzero phase in the Hamiltonian (chirality) hinders excitation transfer between opposite sites of the symmetric cycle (even $N$), as it reduces the detection probability at the target. In contrast, chirality enhances this transfer in the asymmetric cycle (odd $N$) [Fig. 2(b)]. The directionality induced by the nonzero phases causes the detection probability to increase at either of the two target sites $\delta=(N\pm 1)/2$ opposite to the starting one. In particular, for $\tau<\tau^{*}$ the detection probability is always maximized at $\phi_{\rm opt}=\pm\pi/2N$, and minimized at $\phi=0,\pm\pi/N$. Finally, we note that while $P_{\rm det}(\phi,\tau)=P_{\rm det}(-\phi,\tau)$ for even $N$, this symmetry is lifted for odd $N$ [Fig. 2(c,d)]. However, for odd $N$ the detection probability is symmetric upon changing both the target site and the sign of $\phi$.
@@
Optimal detection period $\tau$.—The detection period $\tau$ plays a crucial role, as the detection probability for both even and odd $N$ shows two qualitatively distinct regimes separated by a sharp transition at a critical threshold $\tau^{*}$ [see Fig. 2(a,b)]. When $\tau<\tau^{*}$, $P_{\rm det}(\phi,\tau)$ is smooth, generally increasing with $\tau$ (recall we fixed $T=n\tau$), and remains high over extended intervals. Conversely, when $\tau>\tau^{*}$, a highly structured pattern with sharp oscillations emerges. As $\tau\to 0$, we observe a quantum Zeno effect [45, 46]. Here, frequent measurements performed on a state |$\ket{\delta}$ (orthogonal to the initial one) effectively confine the walker’s evolution in the subspace orthogonal to the detection subspace, preventing it from reaching |$\ket{\delta}$. As a result, the detection probability approaches zero as $\tau\to 0$. A qualitatively analogous behavior is observed when varying the number of sites $N$, with the critical threshold $\tau^{*}$ exhibiting only minor variation. This is illustrated in Fig. 3, where $P_{\rm det}(\phi_{\rm opt},\tau)$ is plotted for different $N$. As evident from Figs. 2-3, although maxima of the detection probability also occur in the region $\tau>\tau^{*}$, attaining them would require a fine tuning of $\tau$ that is difficult to achieve in practice. Since we are interested in a robust protocol, we maximize the detection probability in the relatively wide region $\tau<\tau^{*}$.
@@
Below, we present two complementary interpretations of the results above concerning the optimal parameters $\phi_{\rm opt}$ and $\tau_{\rm opt}$, and the threshold $\tau^{*}$: one based on the analysis of the Perron-Frobenius spectrum, and the other on the analysis of dark states.
Perron-Frobenius analysis.—Local monitoring repeatedly interrupts the unitary evolution of the system by projective measurements, resulting in an overall non-unitary evolution. The elementary step—a free evolution (3) for time $\tau$ followed by the detection attempt $\hat{D}=\outerproduct{\delta}{\delta}$ |—is governed by the (non-unitary) Perron–Frobenius operator  [^1] The authors of Ref. [38] refer to this operator as the survival operator.
@@
$S(n)=\|\hat{O}^{n}(\tau)|\psi_{0}\rangle\|^{2}=\sum_{j=0}^{N-1}|\mu_{j}|^{2n}|\langle\mu_{j}|\psi_{0}\rangle|^{2}$ (6)
@@
is the probability that the detector has not clicked after $n$ detection attempts—that is, the walker has not been detected yet [38]—and relates to the detection probability (4) via $S(n)=1-P_{\rm det}(n)$. The fact that $S(n)\leq 1$ means that $|\mu_{j}|\leq 1\;\forall j$. In particular, the contributions from eigenvectors with $|\mu_{j}|<1$ vanish for $n\to\infty$, while those from eigenvectors with $|\mu_{j}|=1$ survive. The asymptotic dynamics is thus determined by the largest-modulus eigenvalues and their corresponding eigenvectors. The analysis of the eigenproblem of $\hat{O}(\phi,\tau)$ will equip us with the necessary tools to understand both the presence of the threshold $\tau^{*}$ and the optimal values of the parameters. For clarity of discussion, we focus here on the case of odd $N$, deferring the case of even $N$ to Sec. S2 of SM [44].
@@
The existence of the threshold $\tau^{*}$ is reflected in the Perron-Frobenius spectrum—evaluated at fixed $N$ and $\phi=\phi_{\rm opt}$—which differs between the regimes $\tau<\tau^{*}$ or $\tau>\tau^{*}$ [Fig. 4(a,b)]. For $0<\tau<\tau^{*}$, all the eigenvalues fall within the unit circle, $|\mu_{j}|<1\;\forall j$ [see also the largest-modulus eigenvalue in Fig. 4(d)], meaning that all the modes decay and so the excitation will be transferred to the target state with unit probability in the long-time limit, $P_{\rm det}(n\to\infty)=1$. For $\tau>\tau^{*}$, instead, there may exist unit-modulus eigenvalues, indicating the emergence of modes which do not decay over time  [^2] Regardless of $\tau$, $\phi$, and $N$ (even or odd), we always observe the presence of a null eigenvalue with corresponding eigenstate $|\mu=0\rangle=\hat{U}(\phi,-\tau)|\delta\rangle$, which trivially leads to $\hat{O}(\phi,\tau)|\mu=0\rangle\equiv 0$.. These results align with the numerical ones discussed above [Fig. 2(b,d)], including the estimate $\tau^{*}\approx 1.55$. The latter, in turn, represents the upper bound on the detection period, $\tau<\tau^{*}$, over which maximizing the detection probability to have a robust protocol.
@@
The optimal parameters, $\tau_{\rm opt}$ and $\phi_{\rm opt}$, can be estimated by minimizing the Perron-Frobenius eigenvalue $\mu_{\rm PF}$ with the largest modulus, which governs the asymptotic dynamics in the long-time limit. The study of $|\mu_{\rm PF}|$ as a function of $\phi$ and $\tau$ in Fig. 4(c) reveals the following. At $\phi=\phi_{\rm opt}=\pm\pi/2N$ (odd $N$), $|\mu_{\rm PF}|$ decreases smoothly with increasing $\tau<\tau^{*}$, attains a minimum at $\tau_{\rm opt}$, and for $\tau>\tau^{*}$ its behavior becomes highly oscillatory, with $|\mu_{\rm PF}|=1$ for specific values of $\tau$ [Fig. 4(d)]. At a given detection period $\tau<\tau^{*}$, $|\mu_{\rm PF}|$ is minimum at the optimal phase $\phi_{\rm opt}$ [Fig. 4(e)]. This behavior with respect to $\phi$ and $\tau$ is consistent with that of the detection probability in Fig. 2(b,d).
@@
A remark is in order. The Perron-Frobenius analysis characterizes the asymptotic behavior in the long-time limit, whereas previously we assumed a finite observation time $T$, which may not correspond to the long-time limit (we will discuss it at the end of the Letter). In this regard, letting $\tau_{\rm PF}$ be the value for which $|\mu_{\rm PF}|$ is minimum, we observe that $\tau_{\rm PF}$ is well approximated by $\tau_{\rm opt}$ for low $N$, and this agreement extends to higher $N$ upon increasing the total observation time $T$ (data not shown). In other words, the Perron-Frobenius analysis can be used to determine the optimal detection period, provided the total observation time is large enough. Below, we discuss the fact that $\tau_{\rm PF}$ approaches the value $\pi/2$ as $N$ increases  [^3] This trend is not captured by Fig. 3, where the maximum of the detection probability shifts instead toward lower $\tau$, due the finite observation time $T$., in terms of dark states.
@@
Figure 4: Eigenvalues of the Perron-Frobenius operator (Optimal quantum transport on a ring via locally monitored chiral quantum walks) on the unit circle for two representative values of the detection period, (a) $\tau<\tau^{*}$ and (b) $\tau>\tau^{*}$ at $\phi=\pi/2N$, $\delta=10$. (c) Density plot of the largest-modulus PF eigenvalue $|\mu_{\rm PF}|$ as a function of $\phi$ and $\tau$. (d) Curve $|\mu_{\rm PF}|$ as a function of $\tau$ at fixed $\phi=\pi/2N$. (e) Curve $|\mu_{\rm PF}|$ as a function of $\phi$ at fixed $\tau=1.53$. Other parameters are: $N=21$, $\delta=10$.
@@
Dark state analysis.—As thoroughly explained in [38], the presence of a detector in a finite-dimensional quantum system, such as ours, splits the Hilbert space into two parts: a bright subspace, whose states are guaranteed to be eventually detected, ($P_{\rm det}(\infty)=1$), and a dark subspace, whose states are never detected, not even after an infinite number of detection attempts ($P_{\rm det}(\infty)=0$). A state that lies in the dark subspace is orthogonal (and remains orthogonal under the action of the evolution operator) to the detection site $|\delta\rangle$. A state that satisfies these properties—i.e., it is stationary with respect to the evolution and detection attempts—is called a dark state. The total detection probability $P_{\rm det}(\infty)$ is the squared modulus of the projection of the initial state $|\psi_{0}\rangle$ on the bright space [38]. Accordingly, whenever $|\psi_{0}\rangle$ has a nonzero projection on the dark space, then $P_{\rm det}(\infty)<1$, meaning that we are not guaranteed to detect the walker even with infinite measurements. Understanding when dark states arise and characterizing them is therefore important to optimize the detection probability.
@@
Considering the time evolution in Eq. (3), a dark state can be built from the energy eigenstates (2),